Find 𝑑𝑦 𝑑π‘₯ if π‘₯ 2𝑦 + 2𝑦 3 = 3π‘₯ + 2𝑦


1
Expert's answer
2022-04-05T16:38:53-0400

"x\u00b2y+2y\u00b3=3x+2y"


"2xy+x²\frac{dy}{dx}+6y²\frac{dy}{dx}=3+2\frac{dy}{dx}"


"x²\frac{dy}{dx}+6y²\frac{dy}{dx}-2\frac{dy}{dx}=-2xy+3"


"(x²+6y²-2)\frac{dy}{dx}=3-2xy"


"\frac{dy}{dx}=\frac{3-2xy}{x²+6y²-2}"